OpenStudy (anonymous):

Is the line for the equation x = 4 horizontal or vertical? What is the slope of the line? Select the best answer. The line is horizontal. The slope is 0. my choice The line is vertical. The slope is undefined. my second choice The line is horizontal. The slope is undefined. The line is vertical. The slope is 0.

OpenStudy (anonymous):

@sleepyjess when you have time :)

Directrix (directrix):

Associate the x in this question with the word Vertex which ends in x and begins with VER. A line of the form x = number is a vertical line.

OpenStudy (anonymous):

so it would be B?

OpenStudy (anonymous):

sorry, D

Directrix (directrix):

We didn't talk about slope so let's do that. So far, we know the line is vertical.

OpenStudy (anonymous):

the slope would be 0 right

OpenStudy (anonymous):

beause to get to the next point, you rise up 0 and run over 1 to get to your next point, so we have 0/1 = 0

Directrix (directrix):

>>the slope would be 0 right No.

Directrix (directrix):

All points on a vertical line have the same x value. So, when you do the slope formula, you will get zero in the denominator. That is a signal that the slope is undefined.

Directrix (directrix):

>>beause to get to the next point, you rise up 0 and run over 1 to get to your next point, so we have 0/1 = 0 The concept here is correct but .. you never get to run over to the right 1 unit on a vertical line because if you do, then you will be off the vertical line.

OpenStudy (anonymous):

oh sorry was thinking about a horizontal line

Directrix (directrix):

What you wrote >>beause to get to the next point, you rise up 0 and run over 1 to get to your next point, so we have 0/1 = 0 is a good explanation of why the slope of a horizontal line is 0.

OpenStudy (anonymous):

yeah i just noticed I was talking about a horizontal line the whole time

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

ok I understand now

Directrix (directrix):

It is easy to get mixed up on this. That is why when I first learned it, I made up a way to remember it.

OpenStudy (anonymous):

do remember that way? i want to not get mixed up on this cause i usually ddo